# Reversal Blessing: Thinking Backward May Outpace Thinking Forward in Multi-choice Questions

### v2

@@ Abstract
Language models usually use left-to-right (L2R) autoregressive factorization. However, L2R factorization may not always be the best inductive bias. Therefore, we investigate whether alternative factorizations of the text distribution could be beneficial in some tasks. We investigate right-to-left (R2L) training as a compelling alternative, focusing on multiple-choice questions (MCQs) as a test bed for knowledge extraction and reasoning. Through extensive experiments across various model sizes (2B-8B parameters) and training datasets, we find that R2L models can significantly outperform L2R models on several MCQ benchmarks, including logical reasoning, commonsense understanding, and truthfulness assessment tasks. Our analysis reveals that this performance difference may be fundamentally linked to multiple factors including calibration, computability and directional conditional entropy. We ablate the impact of these factors through controlled simulation studies using arithmetic tasks, where the impacting factors can be better disentangled. Our work demonstrates that exploring alternative factorizations of the text distribution can lead to improvements in LLM capabilities and provides theoretical insights into optimal factorization towards approximating human language distribution, and when each reasoning order might be more advantageous.
[…]
Reversal Blessing: Thinking Backward May Outpace Thinking Forward in Multi-choice Questions
Yizhe Zhang1thanks: Equal contribution., Richard He Bai [^1] footnotemark: 1, Zijin Gu1thanks: Core contribution., Ruixiang Zhang1, Jiatao Gu1, Emmanuel Abbe1, Samy Bengio1, Navdeep Jaitly1 1Apple
@@ 1 Introduction
Large Language Model (LLM) pretraining commonly employs left-to-right (L2R) next-token prediction, an approach that enables efficient parallelization and caching. This method models the text distribution $p(x)$ as a factorized autoregressive chain as $p(x_{t}|x_{<t})$. L2R naturally aligns with human cognitive processes of text generation and reasoning, making it well-suited for inference tasks. However, while perfect modeling of each $p(x_{t}|x_{<t})$ would theoretically enable exact recovery of the data distribution $p(x)$, neural networks inevitably introduce approximation errors for each $p(x_{t}|x_{<t})$. These errors compound over timestep $t$ during inference, potentially resulting in hallucinations and repetitions in generation (Bengio et al., 2015; Zhang et al., 2023). Further, L2R factorization can result in inductive biases that lead to unwanted behaviors. For example, Allen-Zhu and Li (2023a) show that inverse search is challenging for L2R LLMs.
@@
We wonder whether L2R is optimal, and if alternative factorizations might capture unique aspects of the data distribution that complement L2R. Can specific factorizations achieve lower approximation errors compared to L2R, or reduce L2R’s inherent bias in particular task domains?
@@
Autoregressive modeling in right-to-left (R2L) fashion factorizes $p(x)$ as $p(x_{t}|x_{>t})$, which presents a particularly promising alternative that has been examined in previous work (Papadopoulos et al., 2024; Berglund et al., 2023; Zhang-Li et al., 2024). This setup views the task as predicting the previous token, and it can achieve prediction losses comparable to the L2R next token prediction objective, due to its symmetry to L2R. We leave the exploration of finding the optimal factorization for language data as future work.
@@
We investigates three questions: (1) How to evaluate R2L models on knowledge extraction and basic reasoning tasks? (2) Can R2L factorization match or surpass L2R’s capabilities in knowledge extraction and reasoning for downstream tasks? (3) What are underlying factors determining the preference of L2R or R2L factorizations?
@@
To address these questions, we conducted controlled experiments comparing L2R and R2L models trained with identical data and computational resources. We evaluated both factorization approaches using standard LLM benchmarks with Multiple-Choice Questions (MCQs). For simplicity, we limit our comparison to MCQs, and leave the evaluations for generative tasks as future work. For R2L models, we applied Bayesian inference to implement "reverse thinking," evaluating choices based on their likelihood of generating the prompt (Figure 1). Our results demonstrate that R2L factorization consistently outperforms L2R across various model sizes and pretraining datasets on several MCQ reasoning and knowledge extraction tasks.
@@
We then investigate the underlying reasons that contribute to R2L’s superior performance in certain scenarios and aim to establish principles for selecting between L2R and R2L factorizations. We hypothesize that the effectiveness of L2R and R2L may be fundamentally linked to several factors: calibration, computability, and the conditional entropy of the factorization direction. Specifically, we found that the factorization direction that achieves lower conditional perplexity generally yields better evaluation results. Nevertheless, these factors are intricately interwoven in actual MCQs, complicating the analysis. To disentangle these factors and ablate on their impact to the performance of L2R or R2L factorization, we design a controlled simulation study using arithmetic tasks, revealing how various factors influence the effectiveness of certain factorization. Our code and model checkpoints have been made available to facilitate future research. [^1] https://github.com/apple/ml-reversal-blessing.
@@ 2.1 Solving MCQs
As shown in Figure 1, in MCQs, LLM process a question $q$ alongside a set of answer choices $A=\{a_{1},a_{2},\ldots,a_{n}\}$. Each (question, answer) pair $(q,a_{i})$ is encoded to compute a relevance score $s_{i}$. The model then selects the answer $a_{k}$ corresponding to the highest score: $k=\arg\max_{i}s_{i}$.
@@
If an R2L model is trained, $s_{i}$ can be computed using Bayes’ rule:
$s_{i}$ $=\log p(a_{i}\mid q)/M_{i}$
$=\frac{1}{M_{i}}(\log p_{R2L}(q\mid a_{i})+\log p_{R2L}(a_{i})-C),$
where $M_{i}=\text{len}(q,a_{i})$, $C=\log p_{R2L}(q)$ is a constant. $\log p_{R2L}(q\mid a_{i})$ and $\log p_{R2L}(a_{i})$ can be autoregressively factorized in R2L manner similar to the forward thinking process in Eq. (1). We consider 3 paradigms of the $s_{i}$ for reverse thinking: (1) normalized $s_{i}$ with $M_{i}=\text{len}(q,a_{i})$ resembling the forward thinking; (2) unnormalized $s_{i}$ with $M_{i}=1$; (3) unnormalized $s_{i}$ without prior, i.e. $s_{i}=\log p_{R2L}(q\mid a_{i})$.
@@ 2.2 Model evaluation
We conduct our evaluation on standard LLM evaluation tasks with MCQs that cover different domains including commonsense reasoning, logical reasoning, truthfulness evaluation and more.
@@ 2.3 Model Pretraining
To pretrain the model, we first tokenize each complete dataset. The R2L model is then trained by reversing all tokens within each training data instance. For a fair comparison between the R2L and L2R models, both models are pretrained from scratch using the same Fineweb-EDU subset dataset comprising 350B tokens (Penedo et al., 2024). Each model consists of 2B parameters (EDU-2B). This is the default setting in our experiments. Both the L2R and R2L models are trained for a single epoch, ensuring each training instance is seen only once, thus the training loss should align with the validation loss. More details for model architecture and training are provided in Appendix A.
To validate the robustness of our findings, we further train two additional variants of settings. These include 8B L2R and R2L models trained with the same 350B Fineweb-EDU dataset (EDU-8B), and 2B L2R and R2L models trained with a random subset of the DCLM dataset (Li et al., 2024a) containing 350B tokens (DCLM-2B).
@@ 2.4 Results
We present our results in Table 1. To verify our pretraining pipeline, we first compare the performance of our pretrained model with the 2B model trained by Huggingface (Penedo et al., 2024) (HF-2B)  [^2] https://huggingface.co/spaces/HuggingFaceFW/blogpost-fineweb-v1. Under similar model size and the same dataset, our 2B model’s performance (EDU-2B) is either comparable to or exceeds the L2R results reported by Huggingface HF-2B.
We then compare the L2R and R2L on all evaluated tasks. As shown in Table 1, surprisingly, for 4 out of the 11 tasks (LogiQA, OpenbookQA, TruthfulQA and Commonsense QA), using R2L with reverse thinking actually improves reasoning performance. In some cases, the improvement was substantial (e.g. 51.23% on TruthfulQA). These results held consistently across different model sizes (2B, 8B), datasets (DCLM, FineWeb EDU), and random seeds, suggesting it is not merely random fluctuation.
@@
For reverse thinking with R2L, we use the paradigm 3 (i.e., unnormalized $s_{i}$ without prior) for downstream tasks evaluation. We compare the three paradigms for reverse thinking in Appendix B, Table 4. Ideally, $s_{i}$ should incorporate priors, as in paradigm 1 or 2. However, in practice, using $s_{i}$ without prior (paradigm 3) consistently yields the best performance except for Social IQA and PIQA. We hypothesize this may be due to intrinsic difficulty of estimating the prior probabilities $p(a)$ using LLMs, due to the "surface competition" calibration issues (Holtzman et al., 2021). We provide detailed explanation of our hypothesize using an illustrative example in Appendix D. Since MCQ answer choices are generally designed to be reasonable text, assigning a uniform prior is probably a reasonable approach.
@@
We also monitor the training loss for pretraining the models on both directions. We observed findings similar to Papadopoulos et al. (2024) in that L2R yields a lower loss compared to R2L, even though both model the same target data distribution. In Papadopoulos et al. (2024), the largest model that was trained had $405$M parameters while our models were trained at the popular small LLM size range of 2B-8B parameters. At this size, we observe a similar percentage difference as reported by previous work, of about 2%-2.5% increase in loss when using R2L, indicating learning the R2L factorization is more challenging. This makes it particularly interesting that on a bunch of MCQ tasks we see the R2L is performing better, as elaborated above.
@@ 3.1 Calibration
The first potential explanation concerns the scoring mechanism in forward thinking, where $s_{i}=\log p_{L2R}(a_{i}\mid q)$. Eq. (1) might not lead to an optimal estimation of $p(a|q)$ as it suffers from several calibration issues. Among the choices, some may contain more words that are highly predictable (e.g., "Hong Kong" or stop-words like "a"), potentially leading to spuriously inflated relevance scores. Additional, Holtzman et al. (2021) shows that simple probability normalization in MCQs is challenging because different surface forms of semantically equivalent answers compete for probability mass, potentially diluting scores for correct answers due to this "surface form competition".
In contrast, reverse thinking with paradigm 3, where $s_{i}=\log p_{R2L}(q\mid a_{i})$, mitigates this issue since the target question $q$ remains constant across all choices. We provide rationale analysis on how R2L paradigm 3 alleviates "surface competition" in Appendix D. In a nutshell, forward thinking suffers from surface form competition, where semantically similar words (e.g., "dog" and "puppy") split probability mass, reducing the likelihood of selecting the correct answer. Reverse thinking mitigates this by enforcing a uniform prior, eliminating competition in the prior distribution and allowing a fairer comparison between answer choices.
This suggests that reverse thinking inherently "auto-normalizes" different choices, resulting in more robust evaluation. However, this sole theory fails to explain why reverse thinking does not consistently outperform forward thinking across all tasks, instead showing superior performance only in specific MCQ scenarios.
@@ 3.2 Computability
It is tempting to consider this computational complexity asymmetry as the main underlying cause for why L2R or R2L is preferred for specific tasks. However, recent research (Mirzadeh et al., 2024; Kambhampati, 2024; Valmeekam et al., 2024) find that LLMs may not actually perform genuine reasoning or computing, as evidenced by their poor generalization when tasks undergo minor modifications. This implies that LLMs mainly emulate reasoning patterns from their training data instead of carrying out actual logical computation, weakening the hypothesis that directional preferences stem from varying computability in different directions.
Furthermore, most MCQs primarily involve knowledge retrieval and basic reasoning, which might not reach the complexity threshold where computational hardness would become a significant factor. Therefore, acknowledging that computability may be a factor, we keep exploring alternative hypotheses.
Figure 3: Lower conditional entropy is typically associated with higher accuracy in the reasoning direction.
@@ 3.3 Conditional Entropy
Previous work (Berglund et al., 2023; Allen-Zhu and Li, 2023b) has also demonstrated that LLMs suffer from the "reversal curse", indicating that inverse R2L search in LLMs is inherently challenging for L2R models - due the disconnect between training and inference directions. Consider an LLM trained on sequences of knowledge/information name entities $(e_{1},e_{2},\cdots,e_{n})$. LLM may effectively construct a directed search graph that maps the key $(e_{1},\cdots,e_{i-1})$ to the value $e_{i}$ for any $i$. Following this logic, the training data essentially forms a Bayesian network that can be represented as a directed acyclic graph (DAG) of entities. Similarly, training an R2L model yields an analogous DAG but with reversed edge directions (see Figure 2 for an illustration). The search efficiency between these two graphs may vary given different queries.
We hypothesize here that between two different factorizations of the data, the direction yielding lower conditional entropy will perform better in MCQs, as it reflects better efficiency in knowledge extraction and multi-hop search. We note however, that this is only true when models under both factorization directions have sufficiently low error, which seems to be true for our models here.
More formally, for a downstream MCQ task $T$ with question and answer choices following task-specific data distribution $P_{T}(q,a)$, we compare the conditional entropy in both directions under pretrained L2R and R2L models (Eq. (2) for L2R and Eq. (3) for R2L):
@@
We assume that the conditional entropy is a proxy for the quality of the learned model, and the direction with lower conditional entropy should perform better. However, computing these summations in (2) and (3) is intractable due to the exponentially large candidate space. Therefore, we employ Monte Carlo estimation of (2) and (3) as proxy measures, specifically computing
@@
To verify this hypothesis, we estimate the conditional entropy for all the evaluation tasks. We provide more experimental details in Appendix C. Figure 3 presents our empirical results, which support this hypothesis that lower conditional entropy is typically linked to greater task accuracy, except for CommonsenseQA and OpenbookQA which are outliers likely because of other confounding factors including the computability.
In Figure 3, we observed that the conditional entropy of R2L is generally greater than L2R. This trend could be related to the findings presented in Table 1, indicating that R2L tends to have higher training loss too. Complementing the rationale in Papadopoulos et al. (2024), we hypothesize that the ease with which the language model can approximate the factorized distribution of L2R and R2L, may be also tied to which direction exhibits higher branching factors in that direction. We leave this exploration for future study.

### v3

# What Makes the Preferred Thinking Direction for LLMs in Multiple-choice Questions?

@@ Abstract
Language models usually use left-to-right (L2R) autoregressive factorization. However, L2R factorization may not always be the best inductive bias. Therefore, we investigate whether alternative factorizations of the text distribution could be beneficial in some tasks. We investigate right-to-left (R2L) training as a compelling alternative, focusing on multiple-choice questions (MCQs) as a test bed for knowledge extraction and reasoning. Through extensive experiments across various model sizes (2B-8B parameters) and training datasets, we find that R2L models can significantly outperform L2R models on several MCQ benchmarks, including logical reasoning, commonsense understanding, and truthfulness assessment tasks. Our analysis reveals that this performance difference may be fundamentally linked to multiple factors including calibration, computability, and directional conditional entropy. We analyze the impact of these factors through controlled simulation studies using arithmetic tasks, where the impacting factors can be better disentangled. Our work demonstrates that exploring alternative factorizations of the text distribution can lead to improvements in LLM capabilities and provides theoretical insights into optimal factorization towards approximating human language distribution, and when each reasoning order might be more advantageous. Our code and checkpoints are released at https://github.com/apple/ml-reversal-blessing
@@ 1 Introduction
Large Language Model (LLM) pretraining commonly employs left-to-right (L2R) next-token prediction, an approach that enables efficient parallelization and caching. This method models the text distribution $p(x)$ as a factorized autoregressive chain as $p(x_{t}|x_{<t})$. L2R naturally aligns with human cognitive processes of text generation and reasoning, making it well-suited for inference tasks.
However, while perfect modeling of each $p(x_{t}|x_{<t})$ would theoretically enable exact recovery of the data distribution $p(x)$, neural networks inevitably introduce approximation errors for each $p(x_{t}|x_{<t})$. These errors compound over timestep $t$ during inference, potentially resulting in hallucinations and repetitions in generation (Bengio et al., 2015; Zhang et al., 2023). Further, L2R factorization can result in inductive biases that lead to unwanted behaviors. For example, Allen-Zhu & Li (2023a) show that inverse search is challenging for L2R LLMs, and Berglund et al. (2023) demonstrate the "reversal curse" where models trained on forward text data struggle with inverse relationships.
@@
We investigate whether L2R is optimal, and if alternative factorizations might capture unique aspects of the data distribution that complement L2R. Can specific factorizations achieve lower approximation errors compared to L2R, or reduce L2R’s inherent bias in particular task domains?
@@
Autoregressive modeling in right-to-left (R2L) fashion factorizes $p(x)$ as $p(x_{t}|x_{>t})$, which presents a particularly promising alternative that has been examined in previous work (Papadopoulos et al., 2024; Berglund et al., 2023; Zhang-Li et al., 2024). This setup views the task as predicting the previous token, and it can achieve prediction losses comparable to the L2R next token prediction objective, due to its symmetry to L2R. While R2L may seem counterintuitive given human language processing patterns, it may enable more efficient knowledge extraction in certain scenarios by aligning with the natural direction of information flow in those cases, and it could provide complementary inductive biases that help with specific reasoning tasks.
@@
We investigate three questions: (1) How to evaluate R2L models on knowledge extraction and basic reasoning tasks? (2) Can R2L factorization match or surpass L2R’s capabilities in knowledge extraction and reasoning for downstream tasks? (3) What are the underlying factors determining the preference of L2R or R2L factorizations?
@@
To address these questions, we conducted controlled experiments comparing L2R and R2L models trained with identical data and computational resources. We evaluated both factorization approaches using standard LLM benchmarks with Multiple-Choice Questions (MCQs). For simplicity, we limit our comparison to MCQs, and leave the evaluations for generative tasks as future work. For R2L models, we applied Bayesian inference to implement "reverse thinking," evaluating choices based on their likelihood of generating the prompt (Figure 1).
Our results reveal a surprising and previously unobserved empirical finding: R2L models can significantly outperform L2R models on several standard MCQ benchmarks, including tasks requiring logical reasoning, commonsense understanding, and truthfulness assessment. This finding—where backward thinking can sometimes yield superior performance—is a phenomenon not previously observed or widely recognized in the current literature, especially at the scale and across the diverse benchmarks we evaluated. This observation challenges the prevailing fundamental assumption that L2R autoregressive factorization is the default and optimal inductive bias for LLMs.
Beyond this empirical discovery, our work introduces new perspectives on analyzing the performance differences between L2R and R2L models by proposing three potential underlying factors: calibration, computability, and, critically, conditional entropy. The role of conditional entropy in explaining the preferred reasoning direction, and generally understanding reasoning machinery, is a novel theoretical insight introduced in this paper. We empirically verify this hypothesis, showing that lower conditional entropy generally correlates with higher accuracy in the reasoning direction for many tasks. Our primary contribution lies not in demonstrating that one factorization is universally superior, but in developing a principled framework for understanding when and why different factorization directions may be preferred.
@@
Nevertheless, these factors are intricately interwoven in actual MCQs, complicating the analysis. To disentangle these factors and ablate on their impact to the performance of L2R or R2L factorization, we design a controlled simulation study using arithmetic tasks, revealing how various factors influence the effectiveness of certain factorization. Our code and model checkpoints have been made publicly available for reproduction and to facilitate future research.  [^1] https://github.com/apple/ml-reversal-blessing.
@@ 2.1 Solving MCQs
As shown in Figure 1, in MCQs, LLMs process a question $q$ alongside a set of answer choices $A=\{a_{1},a_{2},\ldots,a_{n}\}$. Each (question, answer) pair $(q,a_{i})$ is encoded to compute a relevance score $s_{i}$. The model then selects the answer $a_{k}$ corresponding to the highest score: $k=\arg\max_{i}s_{i}$.
@@
With an R2L model, $s_{i}$ can be computed using Bayes’ rule:
$s_{i}$ $=\log p(a_{i}\mid q)/M_{i}=\frac{1}{M_{i}}(\log p_{R2L}(q\mid a_{i})+\log p_{R2L}(a_{i})-C),$
where $M_{i}=\text{len}(q,a_{i})$, $C=\log p_{R2L}(q)$ is a constant. $\log p_{R2L}(q\mid a_{i})$ and $\log p_{R2L}(a_{i})$ can be autoregressively factorized in R2L manner similar to the forward thinking process in Eq. equation 1. We consider 3 paradigms of the $s_{i}$ for reverse thinking: (1) normalized $s_{i}$ with $M_{i}=\text{len}(q,a_{i})$ resembling the forward thinking; (2) unnormalized $s_{i}$ with $M_{i}=1$; (3) unnormalized $s_{i}$ without prior, i.e. $s_{i}=\log p_{R2L}(q\mid a_{i})$.
@@ 2.2 Model evaluation
We conducted our evaluation on standard LLM evaluation tasks with MCQs that cover different domains including commonsense reasoning, logical reasoning, truthfulness evaluation and more.
@@ 2.3 Model Pretraining
To pretrain the model, we first tokenize each complete dataset. The R2L model is then trained by reversing all tokens within each training data instance. For a fair comparison between the R2L and L2R models, both models are pretrained from scratch using the same Fineweb-EDU subset dataset comprising 350B tokens (Penedo et al., 2024). Each model consists of 2B parameters (EDU-2B), which is the default setting in our experiments. We also train 1.5B, 4B, 8B L2R and R2L models with the same 350B Fineweb-EDU dataset, and 2B L2R and R2L models trained with a random subset of the DCLM dataset (Li et al., 2024a) containing 350B tokens (DCLM-2B). Both the L2R and R2L models are trained for a single epoch, ensuring each training instance is seen only once, thus the training loss should align with the validation loss. More details for model architecture and training are provided in Appendix A.
@@ 2.4 Results
We present our results in Table 1. To verify our pretraining pipeline, we first compare the performance of our pretrained model with the 2B model trained by Huggingface (Penedo et al., 2024) (HF-2B)  [^2] https://huggingface.co/spaces/HuggingFaceFW/blogpost-fineweb-v1. Under similar model size and the same dataset, our 2B model (EDU-2B) achieves performance comparable to or exceeding the L2R results reported by Huggingface HF-2B. Full results for all models ranging from 1.5B to 8B parameters are provided in our appendix B.
We compared L2R and R2L model performance across all evaluated tasks, employing bootstrap sampling (5 replicates, each with 80% resampling with replacement) for statistical robustness. As shown in Table 1, R2L models with reverse thinking exhibited significantly better reasoning performance on 4 out of 11 tasks: LogiQA, OpenBookQA, TruthfulQA, and CommonsenseQA. Statistical significance results are presented in Table 6.
These results remained consistent across different model sizes (1.5B to 8B), datasets (DCLM, FineWeb EDU), and random seeds, indicating the findings are not due to random fluctuation. The relative performance gain or loss when switching to R2L remained generally stable as model size increased (see appendix B).
For TruthfulQA specifically, we observed the most significant performance gain with R2L the improvement was substantial (51.23% on DCLM-2B). We hypothesize that the "reverse thinking" may inherently align better with truthfulness assessment, as it evaluates the question based on each answer choice rather than generating answers from the question. This framing might help the R2L model better discern subtle inaccuracies that an L2R model might overlook due to "surface form competition". Additionally, R2L models demonstrate significantly lower conditional entropy on TruthfulQA compared to L2R models, which aligns with our hypothesis that lower conditional entropy is associated with higher task accuracy.
@@
For reverse thinking with R2L, we use the paradigm 3 (i.e., unnormalized $s_{i}$ without prior) for downstream tasks evaluation. We compare the three paradigms for reverse thinking in Appendix C, Table 5. Ideally, $s_{i}$ should incorporate priors, as in paradigm 1 or 2. However, in practice, using $s_{i}$ without prior (paradigm 3) consistently yields the best performance except for Social IQA and PIQA. We hypothesize this may be due to intrinsic difficulty of estimating the prior probabilities $p(a)$ using LLMs, due to the "surface competition" calibration issues (Holtzman et al., 2021). We provide detailed explanation of our hypothesize using an illustrative example in Appendix E.
Intuitively, paradigm 3 which uses $p(q\mid a)$ is sensible. In MCQs, answer choices are typically well-formed and reasonable text, meaning their prior probabilities $p(a)$ are unlikely to vary significantly among choices, assigning a uniform prior is probably a reasonable approach. Consequently, $p(a\mid q)$ and $p(q\mid a)$ tend to be highly correlated. Consider a real-world example from Openbook QA: for the question $(q)$A magnet will stick to", candidate answers $(a_{i})$ include "a belt buckle", "a wooden table", "a plastic cup", and "a paper plate". A model can deduce that "a belt buckle" is far more likely to be associated with the question "A magnet will stick to" compared to the other options, demonstrating how $p(q\mid a)$ can effectively capture the relevance between question and answer.
@@
We also monitor the training loss for pretraining the models on both directions. We observed findings similar to Papadopoulos et al. (2024) in that L2R yields a lower loss compared to R2L, even though both models model the same target data distribution. In Papadopoulos et al. (2024), the largest model that was trained had $405$M parameters while our models were trained at the popular small LLM size range of 2B-8B parameters. At this size, we observe a similar percentage difference as reported by previous work, of about 2%-2.5% increase in loss when using R2L, indicating learning the R2L factorization is more challenging. This makes it particularly interesting that on a bunch of MCQ tasks we see the R2L is performing better, as elaborated above.
@@ 3.1 Calibration
The first potential explanation concerns the scoring mechanism in forward thinking, where $s_{i}=\log p_{L2R}(a_{i}\mid q)$. Eq. equation 1 might not lead to an optimal estimation of $p(a|q)$ as it suffers from several calibration issues. Among the choices, some may contain more words that are highly predictable (e.g., "Hong Kong" or stop-words like "a"), potentially leading to spuriously inflated relevance scores. Additionally, Holtzman et al. (2021) shows that simple probability normalization in MCQs is challenging because different surface forms of semantically equivalent answers compete for probability mass, potentially diluting scores for correct answers due to this "surface form competition".
In contrast, reverse thinking with paradigm 3, where $s_{i}=\log p_{R2L}(q\mid a_{i})$, mitigates this issue since the target question $q$ remains constant across all choices. We provide rationale analysis on how R2L paradigm 3 alleviates "surface competition" in Appendix E. In a nutshell, forward thinking suffers from surface form competition, where semantically similar words (e.g., "dog" and "puppy") split probability mass, reducing the likelihood of selecting the correct answer. Reverse thinking mitigates this by enforcing a uniform prior, eliminating competition in the prior distribution and allowing a fairer comparison between answer choices. This suggests that reverse thinking inherently "auto-normalizes" different choices, resulting in more robust evaluation. However, this sole theory fails to explain why reverse thinking does not consistently outperform forward thinking across all tasks, instead showing superior performance only in specific MCQ scenarios.
@@ 3.2 Computability
It is tempting to consider this computational complexity asymmetry as the main underlying cause for why L2R or R2L is preferred for specific tasks. However, recent research (Mirzadeh et al., 2024; Kambhampati, 2024; Valmeekam et al., 2024) finds that LLMs may not actually perform genuine reasoning or computing, as evidenced by their poor generalization when tasks undergo minor modifications. This implies that LLMs mainly emulate reasoning patterns from their training data instead of carrying out actual logical computation, weakening the hypothesis that directional preferences stem from varying computability in different directions. Furthermore, most MCQs primarily involve knowledge retrieval and basic reasoning, which might not reach the complexity threshold where computational hardness would become a significant factor. Therefore, acknowledging that computability may be a factor, we keep exploring alternative hypotheses.
Figure 3: Lower conditional entropy may associate with higher accuracy in the reasoning direction.
@@ 3.3 Conditional Entropy
Previous work (Berglund et al., 2023; Allen-Zhu & Li, 2023b) has also demonstrated that LLMs suffer from the "reversal curse", indicating that inverse R2L search in LLMs is inherently challenging for L2R models - due the disconnect between training and inference directions. Consider an LLM trained on sequences of knowledge/information name entities $(e_{1},e_{2},\cdots,e_{n})$. LLM may effectively construct a directed search graph that maps the key $(e_{1},\cdots,e_{i-1})$ to the value $e_{i}$ for any $i$. Following this logic, the training data essentially forms a Bayesian network that can be represented as a directed acyclic graph (DAG) of entities. Similarly, training an R2L model yields an analogous DAG but with reversed edge directions (see Figure 2 for an illustration). The search efficiency between these two graphs may vary given different queries.
We hypothesize here that for two different factorizations of the data, the direction yielding lower conditional entropy will perform better in MCQs, as it reflects better efficiency in knowledge extraction and multi-hop search. We note however, that this is only true when models under both factorization directions have sufficiently low error, which seems to be true for our models here.
More formally, for a downstream MCQ task $T$ with question and answer choices following task-specific data distribution $P_{T}(q,a)$, we compare the conditional entropy in both directions under pretrained L2R and R2L models (Eq. equation 2 for L2R and Eq. equation 3 for R2L):
@@
We assume that the conditional entropy is a proxy for the quality of the learned model, and the direction with lower conditional entropy should perform better. However, computing these summations in equation 2 and equation 3 is intractable due to the exponentially large candidate space. Therefore, we employ Monte Carlo estimation of equation 2 and equation 3 as proxy measures, specifically computing
@@
To verify this hypothesis, we estimate the conditional entropy for all the evaluation tasks. We provide more experimental details in Appendix D. Figure 3 presents our empirical results, which support this hypothesis that lower conditional entropy is typically linked to greater task accuracy, except for CommonsenseQA and OpenbookQA which are outliers likely because of other confounding factors including the computability. In Figure 3, we observed that the conditional entropy of R2L is generally greater than L2R. This trend could be related to the findings presented in Table 1, indicating that R2L tends to have higher training loss too. Complementing the rationale in Papadopoulos et al. (2024), we hypothesize that the ease with which the language model can approximate the factorized distribution of L2R and R2L, may be also tied to which direction exhibits higher branching factors in that direction. We leave this exploration for future study.